please help me solve quadratic equations by factoring: 10x^2=9x
Answer by stanbon(75887)

please help me solve quadratic equations by factoring: 10x^2=9x
------
10x^2-9x = 0
x(10x-9) = 0
x = 0 or x = 9/10
========================
Cheers,
Stan H.